OpenStudy (anonymous):

Mr. Trundle used exactly 128 ceramic tiles to cover the floor of a shower. If each tile is a square with 6-inch sides, what is the area of the shower floor?

OpenStudy (anonymous):

help plz @Anikate

OpenStudy (anonymous):

i give medal plz help

OpenStudy (bahrom7893):

128*(6)^2 = 128*36 = 4608 in ^2

OpenStudy (anonymous):

thx

OpenStudy (anikate):

sry